# Credal Learning Theory

### v3

@@ Abstract
Statistical learning theory is the foundation of machine learning, providing theoretical bounds for the risk of models learnt from a (single) training set, assumed to issue from an unknown probability distribution. In actual deployment, however, the data distribution may (and often does) vary, causing domain adaptation/generalization issues. In this paper we lay the foundations for a ‘credal’ theory of learning, using convex sets of probabilities (credal sets) to model the variability in the data-generating distribution. Such credal sets, we argue, may be inferred from a finite sample of training sets. Bounds are derived for the case of finite hypotheses spaces (both assuming realizability or not) as well as infinite model spaces, which directly generalize classical results.
@@ 1 Introduction
Statistical Learning Theory (SLT) considers the problem of predicting an output $y\in\mathcal{Y}$ given an input $x\in\mathcal{X}$ by means of a mapping $h:\mathcal{X}\rightarrow\mathcal{Y}$, $h\in\mathcal{H}$, called model or hypothesis, belonging to model (or hypotheses) space $\mathcal{H}$. The loss function $l:(\mathcal{X}\times\mathcal{Y})\times\mathcal{H}\rightarrow\mathbb{R}$ measures the error committed by a model $h\in\mathcal{H}$. For instance, the zero-one loss is defined as $l((x,y),h)\doteq\mathbbm{I}[y\neq h(x)]$, where $\mathbbm{I}$ denotes the indicator function. This function assigns a zero value to correct predictions and one to incorrect ones. Input-output pairs are usually assumed to be generated i.i.d. by a probability distribution $P^{\star}$, which is unknown. The expected risk – or expected loss – of the model $h$,
$L(h)\equiv L_{P^{\star}}(h)$ $\doteq\mathbb{E}_{P^{\star}}[l((x,y),h)]$
$=\int_{\mathcal{X}\times\mathcal{Y}}l((x,y),h)P^{\star}(\text{d}(x,y)),$
measures the expected value – taken with respect to $P^{\star}$ – of loss $l$. The expected risk minimizer
$h^{\star}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}L(h),$
@@
is any hypothesis in the given model space $\mathcal{H}$ that minimizes the expected risk. Given a training dataset $D=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ whose elements are drawn independent and identically distributed (i.i.d.) from probability distribution $P^{\star}$, the empirical risk of a hypothesis $h$ is the average loss over $D$. The empirical risk minimizer (ERM), i.e., the model $\hat{h}$ one actually learns from the training set $D$, is the one minimizing the empirical risk. Statistical Learning Theory seeks upper bounds on the expected risk $L(\hat{h})$ of the ERM $\hat{h}$, and in turn, on the excess risk, that is, on the difference between $L(\hat{h})$, and the lowest expected risk $L(h^{\star})$. This endeavor is pursued under increasingly more relaxed assumptions about the nature of the hypotheses space $\mathcal{H}$. Two common such assumptions are that either the model space is finite, or that there exists a model with zero expected risk (realizability).
@@
In real-world situations, however, the data distribution may (and often does) vary, causing issues of domain adaptation (DA) or generalization (DG). Domain adaptation and generalization are interrelated yet distinctive concepts in machine learning, as they both deal with the challenges of transferring knowledge across different domains. The main goal of DA is to adapt a machine learning model trained on source domains to perform well on target domains. In opposition, DG aims to train a model that can generalize well to unseen data/domains not available during training. In simple terms, DA works on the assumption that our source and target domains are related to each other, meaning that they somehow follow a similar data generating probability distribution. DG, instead, assumes that the trained model should be able to handle unseen target data (Piva et al., 2023).
@@
Some attempts to derive generalization bounds under more realistic conditions within classical SLT have been made. However, those approaches are characterized by a lack of generalizability, and the use of strong assumptions. A more detailed account of the state of the art and their limitations is discussed Section 2.
@@
Imprecise Probabilities (IPs) can provide a radically different solution to the construction of bounds in learning theory. A hierarchy of formalisms aimed at mathematically modeling the ‘epistemic’ uncertainty induced by sources such as lack of data, missing data or data which is imprecise in nature (Cuzzolin, 2020; Sale et al., 2023a, b), IPs have been successfully employed in the design of neural networks providing both better accuracy and uncertainty quantification to predictions (Sensoy et al., 2018; Manchingal and Cuzzolin, 2022; Manchingal et al., 2023; Caprio et al., 2023a; Lu et al., 2024; Wang et al., 2024). To date, however, they have never been considered as a tool to address the foundational issues of statistical learning theory associated with data drifting.
1.1 Contributions
This paper provides two innovative contributions: (1) the formal definition of a new learning setting in which models are inferred from a (finite) sample of training sets, rather than a single training set, each assumed to have been generated by a single data distribution (as in classical SLT); (2) the derivation of generalization bounds to the expected risk of a model learnt in this new learning setting, under the assumption that the epistemic uncertainty induced by the available training sets can be described by a credal set, i.e., a convex set of (data generating) probability distributions. The overall framework is illustrated in Figure 1.
Generalized upper bounds under credal uncertainty are derived under three increasingly realistic sets of assumptions, mirroring classical statistical learning theory treatment: (i) finite hypotheses spaces with realizability, (ii) finite hypotheses spaces without realizability, and (iii) infinite hypotheses spaces. We show that the corresponding classical results in SLT are special cases of the ones derived in the present paper.
1.2 Structure of the paper
The paper is structured as follows. First (Section 2) we present the existing work in addressing data distribution shifts in learning theory. We then introduce our new learning framework (Section 3). In Section 4 we illustrate the bounds derived under credal uncertainty, and show how classical results can be recovered as special cases. Section 5 concludes and outlines future undertakings. We prove our results in Appendix A, and we provide a synthetic experiment in Appendix B.
@@ 2 Related Work
The standard statistical approach is based on the assumption that test and training data are i.i.d. according to an unknown distribution. This assumption can fall short in real-world applications: as a result, many recent papers have been focusing on the “Out of Distribution” (OOD) generalization problem, also known as domain generalization (DG), to address the discrepancy between test and training distribution(s) (Gulrajani and Lopez-Paz, 2020; Koh et al., 2021). Extensive surveys of existing methods and approaches to DG can be found in Zhou et al. (2022); Wang et al. (2022).
Although several proposals for learning bounds with theoretical guarantees have been made within DA, only a few attempts have been made in the field of DG (Redko et al., 2020; Deshmukh et al., 2019). Most of the theoretical attempts have focused on kernel methods, starting from the seminal work of Blanchard et al. (2011), spanning to a body of later work (see for example Muandet et al. (2013); Deshmukh et al. (2019); Hu et al. (2020)), or are reliant on strong assumptions (e.g. in relation to the similarity of training and testing domains) that are not directly applicable to broader DG approaches. Other work has focused on providing theoretical grounds showing that minimizing the maximum pairwise distance in the source domains (assumed to form a convex hull) leads to a distance decrease between any pair of domains in the convex hull (Albuquerque et al., 2019).
Researchers have also focused on adaptation to new domains over time, treating DG as an online game and the model as a player minimizing the risk associated with introducing new distributions by an adversary at each step (Rosenfeld et al., 2022). Ye et al. (2021) have attempted to relax assumptions and provide more general bounds, focusing on feature distribution characteristics; the authors have introduced terms related to stability and discriminative power to calculate, by an expansion function, the error bound on unseen domains. Causality principles have been leveraged in this sense, for example by Sheth et al. (2022) and Bellot and Bareinboim (2022), which provide distributional robustness guarantees using causal diagrams and source domain data. Researchers have also explored generalization bounds for DG based on the Rademacher complexity, allowing for the approach to be applicable to a broader range of models (Li et al., 2022).
On the empirical analysis side, Gulrajani and Lopez-Paz (2020) provide a comprehensive review of the state of the art. Though a simple ERM was found to outperform other more sophisticated methods (Cha et al., 2021), this approach is criticized for non-generalizability; Izmailov et al. (2018) highlights the importance of searching for flat minima in the training process for improved generalization.
@@
All the aforementioned approaches take a deterministic, and point estimate-like, stance (i.e., assuming single training set) to the derivation of generalization bounds. In this paper, in opposition, we explicitly acknowledge the uncertainty inherent to domain variation in the form of a sample of training sets, each assumed to be generated by a different distribution, and propose a robust and flexible approach representing the resulting epistemic uncertainty via credal sets.
@@ 3 Credal Learning
Let us formalize the notion of learning a model from a collection of (training) sets of data, each issued from a different ‘domain’ characterized by a single, albeit unknown, data generating probability distribution.
Assume that we wish to learn a mapping (model, or hypothesis) $h:\mathcal{X}\rightarrow\mathcal{Y}$ between an input space $\mathcal{X}$ and an output space $\mathcal{Y}$ (where, once again, the mapping $h$ belongs to a hypotheses space $\mathcal{H}$), having as evidence a finite sample of training sets, $D_{1},\ldots,D_{N}$,
$D_{i}=\{(x_{i,1},y_{i,1}),\ldots,(x_{i,n_{i}},y_{i,n_{i}})\}.$
Assume also that the data in each set $D_{i}$ has been generated by a distinct probability distribution $P^{\star}_{i}$.
The question we want to answer is: What sort of guarantees can be derived on the expected risk of a model learnt from such a sample of training sets? How do they relate to classical Probably Approximately Correct (PAC) bounds from statistical learning theory?
@@ 3.1 Objectivist Modeling
While in classical statistical learning theory results are derived assuming no knowledge about the data generating process, the theorems and corollaries in this paper do require some knowledge, although incomplete, of the true distribution. To be more specific, we will posit that, by leveraging the available evidence $D_{1},\ldots,D_{N}$, the agent is able to elicit a credal set – i.e., a closed and convex set of probabilities – that contains the true data generating process $P^{\text{true}}$ for a new training sample $D_{N+1}$, possibly different from $D_{1},\ldots,D_{N}$. As we shall see in Section 4, though, this extra modeling effort allows us to derive stronger results.
There are at least two ways in which such credal set can be derived, that is, via objectivist and subjectivist modeling. In this section, we present the former, and we start by inspecting the frequentist approach to objectivist modeling.
@@ 3.1.1 Epsilon-contamination models
In classical frequentist statistics, given the available dataset, the agent assumes the analytical form of a likelihood $\mathcal{L}$ (not to be confused with the expected loss function, which we denote by a roman letter $L$), e.g. a Normal or a Gamma. As shown by Huber and Ronchetti (2009), though, small perturbations of the specified likelihood can induce substantial differences in the conclusions drawn from the data. A robust frequentist agent is interested in statistical methods that may not be fully optimal in the ideal model, but still lead to reasonable conclusions if the ideal model is only approximately true (Augustin et al., 2014). To account for this, the agent specifies the class of $\epsilon$-contaminated distributions $\mathcal{P}=\{P:P=(1-\epsilon)\mathcal{L}+\epsilon Q\text{, }\forall Q\}$, where $\epsilon$ is some positive quantity in $(0,1)$, and $Q$ is any distribution on $\mathcal{X}\times\mathcal{Y}$. Wasserman and Kadane (1990) show that $\mathcal{P}$ is indeed a (nonempty) credal set. In view of this robust frequentist goal, then, requiring that the true data generating process belongs to $\mathcal{P}$ is a natural assumption.
In our framework in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, one approach to building the desired credal set is to specify $N$ many likelihoods $\{\mathcal{L}_{i}\}_{i=1}^{N}$ and $\epsilon_{i}$-contaminate each of them to obtain $\mathscr{L}_{i}=\{P:P=(1-\epsilon_{i})\mathcal{L}_{i}+\epsilon_{i}Q\text{, }\forall Q\}$, $i\in\{1,\ldots,N\}$. The credal set $\mathcal{P}$ can then be derived by putting $\mathcal{P}=\text{Conv}(\cup_{i=1}^{N}\mathscr{L}_{i})$, where $\text{Conv}(\cdot)$ denotes the convex hull operator. [^1] It is easy to see that set $\mathcal{P}$ built this way is indeed a credal set. This because it is (i) convex by definition, and (ii) closed because it is the union of finitely many closed sets. An immediate consequence of Wasserman and Kadane (1990) and references therein is that $\mathcal{P}=\text{Conv}(\cup_{i=1}^{N}\mathscr{L}_{i})=\{P:P(A)\geq\underline{\mathcal{L}}(A)\text{, }\forall A\subseteq\mathcal{X}\times\mathcal{Y}\}$, where ¯$\underline{\mathcal{L}}(A)=\min_{i\in\{1,\ldots,N\}}(1-\epsilon_{i})\mathcal{L}_{i}(A)$, for all $A\subseteq\mathcal{X}\times\mathcal{Y}$. A simple numerical example for this procedure is given in Appendix C.
We can also derive a credal set in an objectivist way by taking the plausibility function route. To describe it, we first need to introduce some concepts from the Dempster-Shafer theory of evidence (Dempster, 1967; Shafer, 1976).

### v4

@@ Abstract
Statistical learning theory is the foundation of machine learning, providing theoretical bounds for the risk of models learned from a (single) training set, assumed to issue from an unknown probability distribution. In actual deployment, however, the data distribution may (and often does) vary, causing domain adaptation/generalization issues. In this paper we lay the foundations for a ‘credal’ theory of learning, using convex sets of probabilities (credal sets) to model the variability in the data-generating distribution. Such credal sets, we argue, may be inferred from a finite sample of training sets. Bounds are derived for the case of finite hypotheses spaces (both assuming realizability or not), as well as infinite model spaces, which directly generalize classical results.
@@ 1 Introduction
Statistical Learning Theory (SLT) considers the problem of predicting an output $y\in\mathcal{Y}$ given an input $x\in\mathcal{X}$ using a mapping $h:\mathcal{X}\rightarrow\mathcal{Y}$, $h\in\mathcal{H}$, called model or hypothesis, belonging to a model (or hypotheses) space $\mathcal{H}$. The loss function $l:(\mathcal{X}\times\mathcal{Y})\times\mathcal{H}\rightarrow\mathbb{R}$ measures the error committed by a model $h\in\mathcal{H}$. For instance, the zero-one loss is defined as $l((x,y),h)\doteq\mathbbm{I}[y\neq h(x)]$, where $\mathbbm{I}$ denotes the indicator function, and assigns a zero value to correct predictions and one to incorrect ones. Input-output pairs are usually assumed to be generated i.i.d. by a probability distribution $P^{\star}$, which is unknown. The expected risk – or expected loss – of the model $h$, $L(h)\equiv L_{P^{\star}}(h)\doteq\mathbb{E}_{P^{\star}}[l((x,y),h)]=\int_{\mathcal{X}\times\mathcal{Y}}l((x,y),h)P^{\star}(\text{d}(x,y))$, measures the expected value – taken with respect to $P^{\star}$ – of loss $l$. The expected risk minimizer $h^{\star}\in\operatorname*{arg\,min}_{h\in\mathcal{H}}L(h)$ is any hypothesis in the given model space $\mathcal{H}$ that minimizes the expected risk. Given a training dataset $D=\{(x_{1},y_{1}),\ldots,(x_{n},y_{n})\}$ whose elements are drawn independently and identically distributed (i.i.d.) from probability distribution $P^{\star}$, the empirical risk of a hypothesis $h$ is the average loss over $D$. The empirical risk minimizer (ERM), i.e., the model $\hat{h}$ one actually learns from the training set $D$, is the one minimizing the empirical risk [44]. Statistical Learning Theory seeks upper bounds for the expected risk $L(\hat{h})$ of the ERM $\hat{h}$, and in turn, for the excess risk, that is, the difference between $L(\hat{h})$ and the lowest expected risk $L(h^{\star})$. This endeavor is pursued under increasingly more relaxed assumptions about the nature of the hypotheses space $\mathcal{H}$. Two common such assumptions are that either the model space is finite, or that there exists a model with zero expected risk (realizability).
@@
In real-world situations, however, the data distribution may (and often does) vary, causing issues of domain adaptation (DA) [11] or generalization (DG) [59]. Domain adaptation and generalization are interrelated yet distinctive concepts in machine learning, as they both deal with the challenges of transferring knowledge across different domains. The main goal of DA is to adapt a machine learning model trained on source domains to perform well on target domains. In opposition, DG aims to train a model that can generalize well to unseen data/domains not available during training. In simple terms, DA works on the assumption that our source and target domains are related to each other, meaning that they somehow follow a similar data-generating probability distribution. DG, instead, assumes that the trained model should be able to handle unseen target data.
@@
Attempts to derive generalization bounds under more realistic conditions within classical SLT have been made (see Section 2). Those approaches, however, are characterized by a lack of generalizability, and the use of strong assumptions. A more detailed account of the state of the art and their limitations is discussed in Section 2. In opposition to all such proposals, our learning framework leverages Imprecise Probabilities (IPs) to provide a radically different solution to the construction of bounds in learning theory.
A hierarchy of formalisms aimed at mathematically modeling the ‘epistemic’ uncertainty induced by sources such as lack of data, missing data or data which is imprecise in nature [24, 62, 63], IPs have been successfully employed in the design of neural networks providing both better accuracy and uncertainty quantification to predictions [17, 47, 48, 49, 64, 73]. To date, however, they have never been considered as a tool to address the foundational issues of statistical learning theory associated with data drifting.
Contributions. This paper provides two innovative contributions: (1) the formal definition of a new learning setting in which models are inferred from a (finite) sample of training sets (via either objectivist or subjectivist modeling techniques, as explained in Section 3), rather than a single training set, each assumed to have been generated by a single data distribution (as in classical SLT); (2) the derivation of generalization bounds to the expected risk of a model learned in this new learning setting, under the assumption that the epistemic uncertainty induced by the available training sets can be described by a credal set [41], i.e., a convex set of (data generating) probability distributions.
The overall framework is illustrated in Figure 1. Generalized upper bounds under credal uncertainty are derived under three increasingly realistic sets of assumptions, mirroring classical statistical learning theory treatment: (i) finite hypotheses spaces with realizability, (ii) finite hypotheses spaces without realizability, and (iii) infinite hypotheses spaces. We show that the corresponding classical results in SLT are special cases of the ones derived in the present paper.
@@
Paper outline. The paper is structured as follows. First (Section 2) we present the existing work addressing data distribution shifts in learning theory. We then introduce our new learning framework (Section 3). In Section 4 we illustrate the bounds derived under credal uncertainty and show how classical results can be recovered as special cases. Section 5 concludes and outlines future undertakings. We prove our results in Appendix A, and we provide synthetic experiments on our first two main results (Theorems 4.1 and 4.5) in Appendix B.
@@ 2 Related Work
The standard statistical approach to generalization is based on the assumption that test and training data are i.i.d. according to an unknown distribution. This assumption can fall short in real-world applications: as a result, many recent papers have been focusing on the “Out of Distribution" (OOD) generalization problem, also known as domain generalization (DG), to address the discrepancy between test and training distribution(s) [31, 40, 68]. Extensive surveys of existing methods and approaches to DG can be found in Wang et al. [71], Zhou et al. [79]. Although several proposals for learning bounds with theoretical guarantees have been made within DA, only a few attempts have been made in the field of DG [28, 60]. Most theoretical attempts have focused on kernel methods, starting from the seminal work of Blanchard et al. [12], spanning to a body of later work (see for example Deshmukh et al. [28], Hu et al. [33], Muandet et al. [57]). In this line of work, assumptions related to boundedness of kernels and continuity of feature maps and loss function render the approaches not directly applicable to broader scenarios.
Other work has focused on providing theoretical grounds using domain-adversarial learning method; in this approach, the authors use a convex combination of source domains in order to approximate the target distribution leveraging H-divergence [2]. Ye et al. [75] have attempted to relax assumptions to provide more general bounds, focusing on feature distribution characteristics; the authors have introduced terms related to stability and discriminative power to calculate the error bound on unseen domains, through the use of an expansion function. Nonetheless, as the authors acknowledge, practical challenges arise concerning the estimation of the expansion function and the choice of a constraint on the top model to improve convergence.
Researchers have also focused on adaptation to new domains over time, treating DG as an online game and the model as a player minimizing the risk associated with introducing new distributions by an adversary at each step [61]. However, in scenarios where the training distribution is significantly outside the convex hull of training distributions [2], or because of unmet strong convexity loss function assumption [61], they fall short from achieving robust generalization. Causality principles have been leveraged in this sense, for example by Bellot and Bareinboim [10], Sheth et al. [67], to provide distributional robustness guarantees using causal diagrams and source domain data. However, causal approaches for improving model robustness across varying domains pose important challenges including reliance on domain knowledge. Researchers have also explored generalization bounds for DG based on the Rademacher complexity, allowing for the approach to be applicable to a broader range of models [42]. Though this simplification has a number of practical benefits, models trained under covariate shift assumptions might suffer in terms of robustness to other distribution shift types. On the empirical analysis side, Gulrajani and Lopez-Paz [31] have provided a comprehensive review of the state of the art. Though a simple ERM was found to outperform other more sophisticated methods in benchmark experiments [21], this approach has been criticized for its non-generalizability. In this direction, Izmailov et al. [35] have highlighted the importance of searching for flat minima in the training process for improved generalization.
@@
All the aforementioned approaches take a point estimate-like, stance (i.e., assuming a single training set) to the derivation of generalization bounds. In this paper, in opposition, we explicitly acknowledge the uncertainty inherent to domain variation in the form of a sample of training sets, each assumed to be generated by a different distribution, and propose a robust and flexible approach representing the resulting epistemic uncertainty via credal sets. Related works on the computational complexity specific to the use of credal sets are discussed in Appendix E.
@@ 3 Credal Learning
Let us formalize the notion of learning a model from a collection of (training) sets of data, each issued from a different ‘domain’ characterized by a single, albeit unknown, data-generating probability distribution. Assume that we wish to learn a mapping $h:\mathcal{X}\rightarrow\mathcal{Y}$ between an input space $\mathcal{X}$ and an output space $\mathcal{Y}$ (where, once again, the mapping $h$ belongs to a hypotheses space $\mathcal{H}$), having as evidence a finite sample of training sets, $D_{1},\ldots,D_{N}$, $D_{i}=\{(x_{i,1},y_{i,1}),\ldots,(x_{i,n_{i}},y_{i,n_{i}})\}$. Assume also that the data in each set $D_{i}$ has been generated by a distinct probability distribution $P^{\star}_{i}$. The question we want to answer is: What sort of guarantees can be derived on the expected risk of a model learned from such a sample of training sets? How do they relate to classical Probably Approximately Correct (PAC) bounds from statistical learning theory?
@@ 3.1 Objectivist Modeling
While in classical statistical learning theory results are derived assuming no knowledge about the data-generating process, the theorems and corollaries in this paper do require some knowledge, although incomplete, of the true distribution. To be more specific, we will posit that, by leveraging the available evidence $D_{1},\ldots,D_{N}$, the agent is able to elicit a credal set – i.e., a closed and convex set of probabilities – that contains the true data generating process $P^{\text{true}}\equiv P^{\star}_{N+1}$ for a new set of data $D_{N+1}$ (that we call the test set), possibly different from $D_{1},\ldots,D_{N}$. As we shall see in Section 4, though, this extra modeling effort allows us to derive stronger results.
There are at least two ways in which such a credal set can be derived, that is, via either an objectivist or a subjectivist modeling stance. In this section, we present the former. We start by inspecting the frequentist approach to objectivist modeling, considering in particular epsilon-contamination models (Section 3.1.1) and belief functions models (Sections 3.1.2, 3.1.3). A further objectivist model based on fiducial inference [3, 32] is outlined in Appendix D.
@@ 3.1.1 Epsilon-contamination models
In classical frequentist statistics, given the available dataset, the agent assumes the analytical form of a likelihood $\mathcal{L}$ (not to be confused with the expected loss function, which we denote by a Roman letter $L$), e.g., a Normal or a Gamma distribution. As shown by Huber and Ronchetti [34], though, small perturbations of the specified likelihood can induce substantial differences in the conclusions drawn from the data. A robust frequentist agent is thus interested in statistical methods that may not be fully optimal under the ideal ‘true’ likelihood model, but still lead to reasonable conclusions if the ideal model is only approximately true [8].
To account for this, the agent specifies the class of $\epsilon$-contaminated distributions
$\mathcal{P}=\{P:P=(1-\epsilon)\mathcal{L}+\epsilon Q\text{, }\forall Q\},$
where $\epsilon$ is some positive quantity in $(0,1)$, and $Q$ is any distribution on $\mathcal{X}\times\mathcal{Y}$. Wasserman and Kadane [74] show that $\mathcal{P}$ is indeed a (nonempty) credal set. In view of this robust frequentist goal, then, requiring that the true data generating process belongs to $\mathcal{P}$ is a natural assumption.
In our framework, in which a finite sample of $N$ training sets $\{D_{i}\}_{i=1}^{N}$ is available, one approach to building the desired credal set is to specify $N$ many likelihoods $\{\mathcal{L}_{i}\}_{i=1}^{N}$ and $\epsilon_{i}$-contaminate each of them to obtain $\mathscr{L}_{i}=\{P:P=(1-\epsilon_{i})\mathcal{L}_{i}+\epsilon_{i}Q\text{, }\forall Q\}$, $i\in\{1,\ldots,N\}$. The credal set $\mathcal{P}$ can then be derived by setting
$\mathcal{P}=\text{Conv}(\cup_{i=1}^{N}\mathscr{L}_{i}),$
where $\text{Conv}(\cdot)$ denotes the convex hull operator. [^1] It is easy to see that the set $\mathcal{P}$ built this way is indeed a credal set. This is because it is (i) convex by definition, and (ii) closed because it is the union of finitely many closed sets. An immediate consequence of Wasserman and Kadane [74] and references therein is that $\mathcal{P}=\text{Conv}(\cup_{i=1}^{N}\mathscr{L}_{i})=\{P:P(A)\geq\underline{\mathcal{L}}(A)\text{, }\forall A\subseteq\mathcal{X}\times\mathcal{Y}\}$, where ¯$\underline{\mathcal{L}}(A)=\min_{i\in\{1,\ldots,N\}}(1-\epsilon_{i})\mathcal{L}_{i}(A)$, for all $A\subset\mathcal{X}\times\mathcal{Y}$. A simple numerical example for such a procedure is given in Appendix C.